1.  y'''-4y'=0                   ans, y=c1+c2e2x+c3e-2x
  2. y'' + 9y' + 5y = 0 
1
Expert's answer
2021-12-09T18:55:01-0500

Q1.

y'''- 4y' = 0

Let us substitute y by emx

So y' = memx, y'' = m²emx and y''' = m³emx

Therefore , m³emx - 4memx = 0

=> emx (m³ - 4m) = 0

=> m³ - 4m = 0

This is the auxiliary equation. Solving it we get

m(m² - 4) = 0

=> m(m - 2)(m + 2) = 0

=> m = 0, 2 , -2

So the general solution of the given differential equation is

y = c1 + c2e2x + c3e-2x where c1, c2 and c3 are arbitrary constants.

Q2.

y'' + 9y' + 5y = 0

Let us substitute y by emx

So y' = memx, y'' = m²emx

Therefore, m²emx + 9memx + 5emx = 0

=> emx(m² + 9m + 5 ) = 0

=> m² + 9m + 5 = 0

This is the auxiliary equation. Solving it we get

m = "\frac{-9 ± \sqrt{9²-4.1.5}}{2.1}"

=> m = "\frac{-9 ± \sqrt{61}}{2}" = "\frac{-9}{2} ±\frac{ \sqrt{61}}{2}"

So the general solution of the given differential equation is

y = "c_{1} e^{\frac{-9 + \sqrt{61}}{2}x}+c_{2}e^{\frac{-9 - \sqrt{61}}{2}x}"

=> y = "e^{-\frac{9}{2}x}(c_{1}e^{\frac{\sqrt{61}}{2}x}+c_{2}e^{-\frac{\sqrt{61}}{2}x})" where c1 and c2 are arbitrary constants.